# Neutrinos in colliding neutron stars and black holes

### v1

@@ I Introduction
Compact objects such as black holes and neutron stars provide us with a remarkable laboratory to test the laws of physics in extreme environments that cannot be reproduced by scientific experiments on Earth – or, really, anywhere else in the known Universe. Black holes and neutron stars are the end point of the evolution of massive stars ($M\gtrsim 8M_{\odot}$, with $M_{\odot}$ the mass of the Sun), after these stars run out of nuclear fuel to burn and collapse under their own gravitational fields. Neutron stars have expected radii of only $\sim(10-14)\,{\rm km}$, roughly the size of a small city, but masses of $\sim(1-2)M_{\odot}$. Their cores reach densities $\gtrsim 10^{15}\,{\rm g/cm^{3}}$, higher than that of atomic nuclei. We know very little about the behavior of matter at such densities, and that behavior remains an important open question in nuclear physics today (see e.g. [1] for more information on neutron stars and the properties of high-density matter). Black holes, of course, are even denser than neutron stars. Mathematically, they are point masses surrounded by a region from which nothing can escape their gravitational pull [^1] Or a little more accurately, in the language of general relativity, the strong spacetime curvature of the black hole creates a region from which no information can escape. The boundary of that region is what we call the event horizon of the black hole. Black holes of any mass can, in theory, exist. Those formed in the collapse of massive stars range from a few times the mass of the sun to tens of solar masses. For a black hole of mass $M_{\rm BH}$, the event horizon has a size of $\sim(1.5-15)\left(\frac{M_{\rm BH}}{M_{\odot}}\right)\,{\rm km}$ (the exact size depends on how fast the black hole is rotating) [2]. A typical black hole of mass $M_{\rm BH}\sim(5-10)M_{\odot}$ [3, 4] will thus have a size comparable to that of a neutron star. With so much mass within a very small volume, black holes and neutron stars offer us with unparalled opportunities to study the laws of gravity as well as, in the case of neutron stars, the properties of extremely dense matter.
@@
An overview of the collision process for neutron star binaries is shown on Fig. 1. As the cores of the neutron stars merge, the lower mass neutron star is strongly distorted by the gravitational field of its companion. A small part of the neutron star is ejected into the surrounding interstellar medium, while more matter remains bound in an accretion disk around the remnant. The collision itself results in additional ejection of hot, shocked material in the interstellar medium. As for the merging cores, they either collapse to a black hole (for high-mass systems), or form a more massive neutron star (for low-mass systems).This is an extremely fast process: the merger itself lasts only a few milliseconds! The collapse itself can immediately follow the merger (i.e. on milliseconds timescales), or for intermediate mass system be delayed (on basically any possible time scale). In the few seconds following the merger, most of the matter in the disk is more slowly accreted by the black hole or neutron star remnant, while the rest is ejected in the form of disk winds. These various channels for matter ejection might play a surprisingly important role in the formation of many heavy elements observed today. About half of the elements heavier than iron, including most of the gold, platinum, and uranium in the Universe, are expected to be formed through nuclear reactions in neutron-rich environments. The exact production site of these elements remains unknown, but the matter ejected by neutron star mergers is a likely candidate [10, 11]. The radioactive decay of the heavy nuclei formed in the ejected matter also results in the emission of an observable optical/infrared signal days to weeks after the merger, called a kilonova [12, 13]. We will discuss these processes in more detail later in this chapter. Black hole-neutron star mergers can similarly eject material and form black hole-disk systems through the disruption of the neutron star by its black hole companion. This is however only possible for relatively low mass or rapidly spinning black holes; in other systems, the neutron star plunges into the black hole before being disrupted. More detailed discussions of our current understanding of neutron star mergers can be found e.g. in [14, 15, 16, 17, 18].
@@ II.1 Properties of dense matter
The properties of the cold, dense matter at the core of neutron stars is an important open question in nuclear physics today, as it is tightly tied to the unknown strength of the strong nuclear forces applied on nucleons (neutrons and protons) at close separation. Most of the matter inside of a neutron star is, as the name implies, very neutron rich ($>90\%$ neutrons, with some protons and electrons). Repulsive forces between nucleons are what allows the neutron star to support itself against gravitational collapse in that regime. At the highest densities that can be encountered at the center of the most massive neutron stars, particles including strange quarks can additionally be created, and we may even observe a phase transition to a quark-gluon plasma. When or even whether this happens remains unclear. Interestingly, there is a direct connection between the size of a neutron star of a given mass and the pressure within its core. That pressure is itself directly related to the strength of nuclear interactions between nucleons, and the composition of the core [1]. The maximum mass that a neutron star can reach without collapsing to a black hole is impacted by the same nuclear physics unknowns. As a result, measurements of the sizes and masses of neutron stars have become an important tool in observational investigations of nuclear physics at high densities [20, 21, 22, 23]. Observations of neutron stars are in that respect complemented by measurements of the internal structure of heavy nuclei such as lead [24, 25], which is similarly impacted by the forces between multiple nucleons at close separation.
@@
As this chapter focuses mostly on neutrino physics, we will not cover the properties of dense matter in more detail but refer the interested reader to e.g. [1] for mode details. For our purpose, it will be enough to know that measuring neutron star masses and radii can be extremely valuable to improve our understanding of nuclear physics. In the context of neutron star mergers, the size of a neutron star determines how easy it is to disrupt, and thus how much matter is ejected in that disruption and how massive of an accretion disk will be formed around the post-merger remnant. The size of neutron stars also impacts their separation and, consequently, velocity when they collide. The mass of merging neutron stars determines whether they quickly collapse into a black hole, or can survive as a neutron star. Finally, both mass and radius impact the gravitational wave emission of neutron star binaries. While pre-merger gravitational wave emission is independent of any neutrino physics, the properties of the ejected matter and the resulting observable electromagnetic signals are definitely impacted by neutrinos. As a result, our ability to recover information about the properties of dense matter from electromagnetic emission powered by neutron star mergers depends in part on a sufficient understanding of neutrino physics in these systems.
@@ II.2 Formation of heavy elements: r-process nucleosynthesis
$X^{A}_{Z}+n\rightarrow X^{A+1}_{Z}+\gamma.$ (1)
Unstable neutron-rich nuclei formed through these neutron captures then undergo beta decay
@@
creating more stable, less neutron-rich nuclei. The main difference between the two processes is the relative timescales between neutron captures and beta decays: in the s-process, neutron captures are slower than the beta decay timescale of nuclei close to the valley of stability. Nuclei thus remain close to that valley of stability (in terms of their number of protons and neutrons). This is known to happen in low mass stars towards the end of their life. In the r-process, on the other hand, neutron capture is much faster than beta decay, leading to the production of very neutron rich, highly unstable nuclei. If enough neutrons are present, neutron capture can eventually create heavy elements that are unstable to fission. The fission fragments can then capture more neutrons, continuing the r-process. The r-process is the only known pathway to naturally synthesize the heaviest elements found in nature, the actinides (e.g. U, Pu). The r-process naturally requires very neutron rich material… like the matter ejected during and after neutron star mergers. While signs of r-process nucleosynthesis have been observed in the aftermath of at least one neutron star merger [10, 11], we do not at this point know whether neutron star mergers are sufficiently frequent and eject enough mass to explain the abundances of r-process elements observed in the Universe, or if another production site for the r-process is necessary. For recent reviews of the r-process and potential astrophysical production sites, see e.g. [26, 27]
Figure 2: Abundances of elements synthesized in outflows from accretion disks produced in black hole-neutron star mergers. The dots show the inferred abundances of r-process elements in the solar system. The lines correspond to systems with different disk masses, which lead to different outflow compositions: more neutron-rich outflows produce more heavy elements (red curve, average $Y_{e}=0.28$); less neutron rich outflows produce only lower mass elements (blue curve, average $Y_{e}=0.33$). Figure reproduced from [28]
@@
As the fluid has no net charge, we have $n_{e^{-}}-n_{e^{+}}=n_{p}$ at least in regions of the remnant where most particles are $(p,n,e^{+},e^{-})$. In those regions, $Y_{e}=Y_{p}=(1-Y_{n})$, with $Y_{p,n}$ the proton/neutron fraction. A low electron fraction thus corresponds to neutron-rich matter, and a high electron fraction to proton-rich matter. For the typical properties of merger outflows, the formation of the heavier r-process elements ($A\gtrsim 140$) requires $Y_{e}\lesssim 0.25$[30]. However, in such neutron-rich ejecta, the lower mass r-process elements are underproduced compared with astrophysical and solar system observations. Moderately neutron rich matter ($Y_{e}\sim 0.25-0.5$) on the other hand does not produce the higher mass elements that we observe in nature, but does produce the lower-mass r-process elements. If neutron star mergers are the main source of r-process elements in the Universe, we thus need both very neutron rich outflows and moderately neutron rich outflows. As the matter from the neutron star originally is $>90\%$ neutrons ($Y_{e}<0.1$), this is only possible if the outflows have significant interaction with neutrinos, making them less neutron rich. We should also note that the exact boundary between those two regimes vary with the properties of the outflows (temperature, velocity) and that smaller variations in the electron fraction of the outflows have a non-negligible impact on the exact abundance pattern of the elements created through r-process nucleosynthesis. This is particularly true for the abundance of actinides, which can vary significantly with the neutron-richness of the outflows, and is also observed to be quite variable in observations of low-metallicity stars [31] – i.e. stars formed so early in the evolution of their galaxies that their r-process elements are likely produced in a single event.
@@ II.3 Kilonovae
Observations of kilonovae could play an important role in our understanding of the properties of dense matter, and of the outcome of r-process nucleosynthesis in the ejected matter. The kilonova signal depends on the mass, temperature, composition, and geometry of the ejected matter [19]. Its composition dependence is particularly interesting, due to its connection to r-process nucleosynthesis. We have already mentioned that more neutron rich outflows produced heavier r-process elements. An important consequence of this result is that only outflows with $Y_{e}\lesssim 0.25$ produce lanthanides ($57\leq Z\leq 71$) and actinides ($89\leq Z\leq 102$). These elements are of particular importance because of their many bound-bound transitions at energies within the infrared/optical bands. Their presence in the outlfows thus delays the transition of the matter from optically thick to photons to optically thin. This leads to the kilonova signal becoming visible at a later time – and typically at lower luminosity and with photons of higher wavelength. In the presence of lanthanides/actinides, kilonovae peak in the infrared on a $\sim$ week time scale instead of the optical on a $\sim$ day time scale. Pratically, this means that the observation of the color and duration of a kilonova provides us with very useful information about the outcome of r-process nucleosynthesis in the ejected matter. Because the mass and size of the original neutron stars impact the properties of the ejected matter, kilonovae can also in theory complement gravitational wave observations in our quest to constrain the properties of neutron stars. This is a difficult process, however. To recover that information, one needs a reliable map between the properties of the initial neutron stars and the eventual kilonova signal. This requires detailed predictions of the outflow properties, an understanding of the outcome of r-process nucleosynthesis, and accurate predictions for the opacity of the outflows to photons. Such a map cannot be reliably produced at this time due in part to the difficulty of properly accounting for neutrino-matter interactions in merger simulations. Other important sources of uncertainty include our poor understanding of the properties of the very neutron-rich nuclei involved in the r-process (and of the reaction rates involved in the r-process), the limited accuracy of our predictions for the opacity of lanthanides and actinides to optical/infrared photons, as well as a limited ability to capture important small scale effects in the evolution of highly magnetized fluids in merger simulations. A significant amount of work is thus still required in order to be able to reliably use kilonovae to estimate the parameters of merging neutron stars, with neutrino-matter interactions playing a large role in this process.
@@ III.1 Important reactions
At the temperatures and densities found in neutron star mergers, neutrinos mostly come into play through four broad classes of processes: charged-current reactions, pair production/annihilation, scatterings, and collective flavor oscillations. In the following sections, we provide a brief overview of how each reaction matters in the merger context. We note that in order of magnitude estimates provided for the neutrino mean-free paths, we completely ignore blocking factors and the finite mass of the electrons. These estimates are thus very inaccurate in the highest density regions and in regions with temperature of $\sim 1\,{\rm MeV}$ or lower. They are however reasonable in the regions most important for the qualitative description of neutrino physics in mergers presented below. We refer the reader to [32, 33, 34] for more accurate calculations of reaction rates and to [35, 36] for more detailed reviews of these processes.
@@ III.1.1 Charged current reactions
Figure 3: Temperature of the fluid and location of the surfaces where neutrino decouple from the fluid on a vertical line along the axis of rotation of a post-merger neutron star remnant, $10\,{\rm ms}$ after merger. For each species of neutrinos, two surfaces are shown: farther left is the surface where the optical depth to absorption is $\sim 2/3$, and farther right the surface where the optical depth to scattering is $\sim 2/3$. The two surfaces are nearly the same for $\nu_{e}$ neutrinos, very close for $\bar{\nu}_{e}$ (labeled as $\nu_{a}$ on the figure), and widely separated for $\nu_{x}$. Left of the first surface, neutrinos are in equilibrium with the fluid. In between the two surfaces, neutrino random-walk through the fluid, on average slowly diffusing outward, without being absorbed. Right of the second surface, neutrinos are largely free-streaming. Due to the different temperatures on the decoupling surfaces, the energy of $\nu_{x}$ outside the neutron star is larger than that of $\bar{\nu}_{e}$, which is itself higher than that of $\nu_{e}$. Figure reproduced from [38].
Within a neutron star remnant, we thus transition from a “trapped” region close to the core to a “free-streaming” region on the surface. Within a post-merger disk, typical initial densities are $10^{-4}-10^{-2}$ of the nuclear saturation density and temperature $\sim(1-10)\,{\rm MeV}$. At their highest density, post-merger accretion disks have scale heights of $O(10\,{\rm km})$ and their optical depth to electron-type neutrinos is thus initially up to $O(1)$. Electron neutrinos are not trapped in the disk, but they have a significant probability of being absorbed by the fluid before escaping the disk. Muon and tau neutrinos are more rarely absorbed at those lower densities and temperature.

### v2

@@ I Introduction
Compact objects such as black holes and neutron stars provide us with a remarkable laboratory to test the laws of physics in extreme environments that cannot be reproduced by scientific experiments on Earth – or, really, anywhere else in the known Universe. Black holes and neutron stars are the end point of the evolution of massive stars ($M\gtrsim 8M_{\odot}$, with $M_{\odot}$ the mass of the Sun), after these stars run out of nuclear fuel to burn and collapse under their own gravitational fields. Neutron stars have expected radii of only $\sim(10-14)\,{\rm km}$, roughly the size of a small city, but masses of $\sim(1-2)M_{\odot}$. Their cores reach densities $\gtrsim 10^{15}\,{\rm g/cm^{3}}$, higher than that of atomic nuclei. We know very little about the behavior of matter at such densities, and that behavior remains an important open question in nuclear physics today (see e.g. [1] for more information on neutron stars and the properties of high-density matter). Black holes, of course, are even denser than neutron stars. Mathematically, they are point masses surrounded by a region from which nothing can escape their gravitational pull [^1] Or a little more accurately, in the language of general relativity, the strong spacetime curvature of the black hole creates a region from which no information can escape. The boundary of that region is what we call the event horizon of the black hole. Black holes of any mass can, in theory, exist. Those formed in the collapse of massive stars range from a few times the mass of the sun to tens of solar masses. For a black hole of mass $M_{\rm BH}$, the event horizon has a size of $\sim(1.5-3)\left(\frac{M_{\rm BH}}{M_{\odot}}\right)\,{\rm km}$ (the exact size depends on how fast the black hole is rotating) [2]. A typical black hole of mass $M_{\rm BH}\sim(5-10)M_{\odot}$ [3, 4] will thus have a size comparable to that of a neutron star. With so much mass within a very small volume, black holes and neutron stars offer us unparalled opportunities to study the laws of gravity as well as, in the case of neutron stars, the properties of extremely dense matter.
@@
An overview of the collision process for neutron star binaries is shown on Fig. 1. As the cores of the neutron stars merge, the lower mass neutron star is strongly distorted by the gravitational field of its companion. A small part of the neutron star is ejected into the surrounding interstellar medium, while more matter remains bound in an accretion disk around the remnant. The collision itself results in additional ejection of hot, shocked material in the interstellar medium. As for the merging cores, they either collapse to a black hole (for high-mass systems), or form a more massive neutron star (for low-mass systems).This is an extremely fast process: the merger itself lasts only a few milliseconds! The collapse of the neutron star to a black hole can immediately follow the merger (i.e. on milliseconds timescales), or for intermediate mass system be delayed (on basically any possible time scale). In the few seconds following the merger, most of the matter in the disk is more slowly accreted by the black hole or neutron star remnant, while the rest is ejected in the form of disk winds. These various channels for matter ejection might play a surprisingly important role in the formation of many heavy elements observed today. About half of the elements heavier than iron, including most of the gold, platinum, and uranium in the Universe, are expected to be formed through nuclear reactions in neutron-rich environments. The exact production site of these elements remains unknown, but the matter ejected by neutron star mergers is a likely candidate [10, 11]. The radioactive decay of the heavy nuclei formed in the ejected matter also results in the emission of an observable optical/infrared signal days to weeks after the merger, called a kilonova [12, 13]. We will discuss these processes in more detail later in this chapter. Black hole-neutron star mergers can similarly eject material and form black hole-disk systems through the disruption of the neutron star by its black hole companion. This is however only possible for relatively low mass or rapidly spinning black holes; in other systems, the neutron star plunges into the black hole before being disrupted. More detailed discussions of our current understanding of neutron star mergers can be found e.g. in [14, 15, 16, 17, 18].
@@ II.1 Properties of dense matter
The properties of the cold, dense matter at the core of neutron stars is an important open question in nuclear physics today, as it is tightly tied to the unknown strength of the strong nuclear forces applied on nucleons (neutrons and protons) at close separation. Most of the matter inside of a neutron star is, as the name implies, very neutron rich ($>90\%$ neutrons, with some protons and electrons). Repulsive forces between nucleons are what allows the neutron star to support itself against gravitational collapse in that regime. At the highest densities that can be encountered at the center of the most massive neutron stars, particles including strange quarks can additionally be created, and we may even observe a transition to an exotic phase [^2] For example a transition from standard nuclear matter with quarks bound in hadrons to deconfined quarks, either in a homogeneous superconducting phase [20] or in inhomegeneous phases [21]. When or even whether this happens remains unclear. Interestingly, there is a direct connection between the size of a neutron star of a given mass and the pressure within its core. That pressure is itself directly related to the strength of nuclear interactions between nucleons, and to the composition of the core [1]. The maximum mass that a neutron star can reach without collapsing to a black hole is impacted by the same nuclear physics unknowns. As a result, measurements of the sizes and masses of neutron stars have become an important tool in observational investigations of nuclear physics at high densities [22, 23, 24, 25]. Observations of neutron stars are in that respect complemented by measurements of the internal structure of heavy nuclei such as lead in laboratories on Earth [26, 27], which is similarly impacted by the forces between multiple nucleons at close separation.
@@
As this chapter focuses mostly on neutrino physics, we will not cover the properties of dense matter in more detail but refer the interested reader to e.g. [1] for more details. For our purpose, it will be enough to know that measuring neutron star masses and radii can be extremely valuable to improve our understanding of nuclear physics. In the context of neutron star mergers, the size of a neutron star determines how easy it is to disrupt, and thus how much matter is ejected in that disruption and how massive of an accretion disk will be formed around the post-merger remnant. The size of neutron stars also impacts their separation and, consequently, velocity when they collide. The mass of merging neutron stars determines whether they quickly collapse into a black hole, or can survive as a neutron star. Finally, both mass and radius impact the gravitational wave emission of neutron star binaries. While pre-merger gravitational wave emission is independent of any neutrino physics, the properties of the ejected matter and the resulting observable electromagnetic signals are definitely impacted by neutrinos. As a result, our ability to recover information about the properties of dense matter from electromagnetic emission powered by neutron star mergers depends in part on a sufficient understanding of neutrino physics in these systems.
@@ II.2 Formation of heavy elements: r-process nucleosynthesis
$X^{A}_{Z}+n\rightarrow X^{A+1}_{Z}+\gamma$ (1)
where $X^{A}_{Z}$ denotes a nucleus with $Z$ protons and $(A-Z)$ neutrons. Unstable neutron-rich nuclei formed through these neutron captures then undergo beta decay
@@
creating more stable, less neutron-rich nuclei. The main difference between the r-process and s-process is the relative timescales between neutron captures and beta decays: in the s-process, neutron captures are slower than the beta decay timescale of nuclei close to the valley of stability. Nuclei thus remain close to that valley of stability (in terms of their number of protons and neutrons). This is known to happen in low mass stars towards the end of their life. In the r-process, on the other hand, neutron capture is much faster than beta decay, leading to the production of very neutron rich, highly unstable nuclei. If enough neutrons are present, neutron capture can eventually create heavy elements that are unstable to fission. The fission fragments can then capture more neutrons, continuing the r-process. The r-process is the only known pathway to naturally synthesize the heaviest elements found in nature, the actinides (e.g. U, Pu). The r-process naturally requires very neutron rich material like the matter ejected during and after neutron star mergers. While signs of r-process nucleosynthesis have been observed in the aftermath of at least one neutron star merger [10, 11], we do not at this point know whether neutron star mergers are sufficiently frequent and eject enough mass to explain the abundances of r-process elements observed in the Universe, or if another production site for the r-process is necessary. For recent reviews of the r-process and potential astrophysical production sites, see e.g. [28, 29]
Figure 2: Relative abundances of elements synthesized in outflows from accretion disks produced in black hole-neutron star mergers. The dots show the inferred abundances of r-process elements in the solar system. The lines correspond to systems with different disk masses, which lead to different outflow compositions (parametrized by the electron fraction $Y_{e}$, as defined in Eq. 3): more neutron-rich outflows produce more heavy elements (red curve, average $Y_{e}=0.28$); less neutron rich outflows produce only lower mass elements (blue curve, average $Y_{e}=0.33$). Figure reproduced from [30]
@@
As the fluid has no net charge, we have $n_{e^{-}}-n_{e^{+}}=n_{p}$ at least in regions of the remnant where most particles are $(p,n,e^{+},e^{-})$. In those regions, $Y_{e}=Y_{p}=(1-Y_{n})$, with $Y_{p,n}$ the proton/neutron fraction, defined in the same was as $Y_{e}$ but replacing the numerator $(n_{e^{-}}-n_{e^{+}})$ with $n_{p,n}$. A low electron fraction thus corresponds to neutron-rich matter, and a high electron fraction to proton-rich matter. For the typical properties of merger outflows, the formation of the heavier r-process elements ($A\gtrsim 140$) requires $Y_{e}\lesssim 0.25$[32]. However, in such neutron-rich ejecta, the lower mass r-process elements are underproduced compared with astrophysical and solar system observations. Moderately neutron rich matter ($Y_{e}\sim 0.25-0.5$) on the other hand does not produce the higher mass elements that we observe in nature, but does produce the lower-mass r-process elements. If neutron star mergers are the main source of r-process elements in the Universe, we thus need both very neutron rich outflows and moderately neutron rich outflows. As the matter from the neutron star originally is $>90\%$ neutrons ($Y_{e}<0.1$), this is only possible if the outflows have significant interaction with neutrinos, making them less neutron rich. We should also note that the exact boundary between those two regimes vary with the properties of the outflows (temperature, velocity) and that smaller variations in the electron fraction of the outflows have a non-negligible impact on the exact abundance pattern of the elements created through r-process nucleosynthesis. This is particularly true for the abundance of actinides, which can vary significantly with the neutron-richness of the outflows, and is also observed to be quite variable in observations of low-metallicity stars [33] – i.e. stars formed so early in the evolution of their galaxies that their r-process elements are likely produced in a single event.
@@ II.3 Kilonovae
Observations of kilonovae could play an important role in our understanding of the properties of dense matter, and of the outcome of r-process nucleosynthesis in the ejected matter. The kilonova signal depends on the mass, temperature, composition, and geometry of the ejected matter [19]. Its composition dependence is particularly interesting, due to its connection to r-process nucleosynthesis. We have already mentioned that more neutron rich outflows produce heavier r-process elements. An important consequence of this result is that only outflows with $Y_{e}\lesssim 0.25$ produce lanthanides ($57\leq Z\leq 71$) and actinides ($89\leq Z\leq 102$). These elements are of particular relevance because of their many bound-bound transitions at energies within the infrared/optical bands. Their presence in the outlfows thus delays the transition of the matter from optically thick to photons to optically thin. This leads to the kilonova signal becoming visible at a later time – and typically at lower luminosity and with photons of higher wavelength. In the presence of lanthanides/actinides, kilonovae peak in the infrared on a $\sim$ week time scale instead of the optical on a $\sim$ day time scale. Pratically, this means that the observation of the color and duration of a kilonova provides us with very useful information about the outcome of r-process nucleosynthesis in the ejected matter. Because the mass and size of the original neutron stars impact the properties of the ejected matter, kilonovae can also in theory complement gravitational wave observations in our quest to constrain the properties of neutron stars. This is a difficult process, however. To recover that information, one needs a reliable map between the properties of the initial neutron stars and the eventual kilonova signal. This requires detailed predictions of the outflow properties, an understanding of the outcome of r-process nucleosynthesis, and accurate predictions for the opacity of the outflows to photons. Such a map cannot be reliably produced at this time due in part to the difficulty of properly accounting for neutrino-matter interactions in merger simulations. Other important sources of uncertainty include our poor understanding of the properties of the very neutron-rich nuclei involved in the r-process (and of the reaction rates involved in the r-process), the limited accuracy of our predictions for the opacity of lanthanides and actinides to optical/infrared photons, as well as a limited ability to capture important small scale effects in the evolution of highly magnetized fluids in merger simulations. A significant amount of work is thus still required in order to be able to reliably use kilonovae to estimate the parameters of merging neutron stars, with neutrino-matter interactions playing a large role in this process.
@@ III.1 Important reactions
At the temperatures and densities found in neutron star mergers, neutrinos mostly come into play through four broad classes of processes: charged-current reactions, pair production/annihilation, scatterings, and collective flavor oscillations. In the following sections, we provide a brief overview of how each reaction matters in the merger context. We note that in order of magnitude estimates provided for the neutrino mean-free paths, we completely ignore blocking factors (i.e. corrections to the interaction rate due to the Pauli exclusion principle, which prevents two fermions from being in the exact same quantum state) and the finite mass of the electrons. These estimates are thus very inaccurate in the highest density regions and in regions with temperature of $\sim 1\,{\rm MeV}$ or lower. They are however reasonable in the regions most important for the qualitative description of neutrino physics in mergers presented below. We refer the reader to [34, 35, 36] for more accurate calculations of reaction rates and to [37, 38] for more detailed reviews of these processes.
@@ III.1.1 Charged current reactions
Figure 3: Temperature of the fluid and location of the surfaces where neutrino decouple from the fluid on a vertical line along the axis of rotation of a post-merger neutron star remnant, $10\,{\rm ms}$ after merger. For each species of neutrinos, two surfaces are shown: farther left is the surface where the optical depth to absorption is $\sim 2/3$, and farther right the surface where the optical depth to scattering is $\sim 2/3$. The two surfaces are nearly the same for $\nu_{e}$ neutrinos, very close for $\bar{\nu}_{e}$ (labeled as $\nu_{a}$ on the figure), and widely separated for $\nu_{x}$. Left of the first surface, neutrinos are in equilibrium with the fluid. In between the two surfaces, neutrinos random-walk through the fluid, on average slowly diffusing outward, without being absorbed. Right of the second surface, neutrinos are largely free-streaming. Due to the different temperatures on the decoupling surfaces, the energy of $\nu_{x}$ outside the neutron star is larger than that of $\bar{\nu}_{e}$, which is itself higher than that of $\nu_{e}$. Figure reproduced from [40].
Within a neutron star remnant, we thus transition from a “trapped” region close to the core to a “free-streaming” region on the surface. Within a post-merger disk, typical initial densities are $10^{-4}-10^{-2}$ of the nuclear saturation density and temperature $\sim(1-10)\,{\rm MeV}$. At their highest density, post-merger accretion disks have scale heights of $O(10\,{\rm km})$ and their optical depth to electron-type neutrinos is thus initially up to $O(1)$. Electron neutrinos are not trapped in the disk, but they have a significant probability of being absorbed by the fluid before escaping the disk. Muon and tau neutrinos are more rarely absorbed at those lower densities and temperature, as the neutrinos propagating through the disk do not have enough energy to produce muon and tau leptons.